OpenStudy (anonymous):

Choose the correct simplification of the expression (2x7y)2(y5)3. 4x7y10 4x9y10 4x14y16 4x14y17

OpenStudy (anonymous):

@Luis_Rivera

OpenStudy (anonymous):

It will expand to give: \[\large 4x^{14}y^2 \cdot y^{15}\]

OpenStudy (anonymous):

Now use : \[\large \color{green}{x^a \cdot x^b = x^{a+b}}\]

OpenStudy (anonymous):

You are only one step away from your answer.. Look carefully, and then use the above formula that I gave in green..